Two closed vessel A and B of equal volume of 8.21L are connected by a narrow tube of negligible volume with open valve.The left hand side container is found to contain 3 mole CO2 and 2 mole He at 400 K.What is the partial pressure of He in vessel B at 500K
 

Grade:

1 Answers

Arun
5 years ago

P(CO2)=3*0.0821*400/8.21=12atm

 

P(He)=2*0.0821*400/8.21=8atm

 Pressure at 55k

 

P1/T1=P2/T2

 

T1=400

T2=500

P1=8atm

8/400=p2/500

P2=4000/400=10atm.=the partial pressure of He in vessel B at 500 K